1. Calculate the first four moments about the mean from the following frequency distribution. 

      Also calculate the Moment Ratios b1 and b2.                         

Groups

2------4

4-------6

6------8

8-----10

10----12

12----14

14----16

16----18

18----20


Frequency

18

24

47

80

102

66

40

21

15


                              


1
Expert's answer
2021-06-01T09:07:14-0400

Given data is:






Fist four moments about mean:

"\mu_1'=\dfrac{1}{n}\sum x_if_i=\dfrac{4445}{413}=10.762"


"\mu_2'=\dfrac{1}{n}\sum x_i^2f_i=\dfrac{48110}{413}=116.489"


"\mu_3'=\dfrac{1}{n}\sum x_i^3f_i=\dfrac{699749}{413}=1694.30"


"\mu_4'=\dfrac{1}{n}\sum x_i^4f_i=\dfrac{9775376}{413}=23669.19"



Now,

"\mu_1=\mu_1'=10.762\\\ \\\mu_2=\mu_2'-\mu_1'=105.727\\\ \\\mu_3=\mu_3'-3\mu_2'\mu_1'+2\mu_1'^3=1694.30-3760.96+2492.92=426.26"


"\mu_4=\mu_4'-4\mu_3'\mu_1'+6\mu_2'\mu_1'^2-3\mu_1'^4=23669.19-72936.2264+80950.98-3739.38\\\mu_4=27944.563"


Now,


"b_1=\dfrac{\mu_3^2}{\mu_2^3}=\dfrac{(426.26)^2}{(105.727)^3}=\dfrac{181697.58}{1181837.39}=0.1537"



"b_2=\dfrac{\mu_4}{\mu_2^2}=\dfrac{27944.563}{11178.19853}=2.499"